Through (-3,5) and (5,3). Thanks
Answer by tutorcecilia(2152)

(-3,5) and (5,3) [Find the slope of the line]
(y-y2)/x-x2) [Formula for a slope]
(3-5)/(5-(-3)) [Plug-in the values]
-2/8=-1/4 [Slope of the line]
.
y=mx+b [Use the slope-intercept form of a line]
5=(-1/4)(-3)+b [Plug-in the values for either of the points]
5=3/4+b [Solve for b]
5-3/4=b
(20-3)/4=b
17/4=b
.
y=mx+b [Plug-in the value of the slope and the y-intercept (b)]
y=-1/4x+17/4 [The equation of the line]